OpenStudy (anonymous):

How do I solve this equation? 1. (5/2x) - (2/3) = (1/x) + (5/6) *I will post what I did so far.

OpenStudy (anonymous):

is it \(\frac{5}{2x}\) or \(\frac{5}{2}x\) ?

OpenStudy (anonymous):

The first one.

OpenStudy (anonymous):

\[\frac{ 5 }{ 2x }\] sorry for any confusion.

OpenStudy (anonymous):

\[\frac{ 15 - 4x}{ 6x } = \frac{ 6 + 5x }{ 6x }\] Since they both have the same denominators, when you cross multiply, they will cancel out anyways so I then did: 15 - 4x = 6 + 5x 15 - 9x = 6 -9x = -10 \[x=\frac{ 10 }{ 9 }\] I plugged the answer in for x and I did not get the same answer for the other side.

OpenStudy (anonymous):

i would add \(\frac{2}{3}\) to both sides, subtract \(\frac{1}{x}\) from both sides, and then cross multiply

OpenStudy (anonymous):

maybe that method would work as well, but \(6-15=-9\) and not \(-10\)

OpenStudy (anonymous):

Whoops, well that's embarrassing :o lol. Thank you very much for your help!

OpenStudy (anonymous):

ye